OpenStudy (anonymous):

What are fibonacci numbers? What is the golden ratio? How do fibonacci numbers and the golden ratios relate to each other?

OpenStudy (anonymous):

Also, where does the golden ratio appear?

OpenStudy (turingtest):

The Fibonacci sequence is a sequence generated by adding the number before itself: 1,1,2,3,5,8,13,21,... because 1+0=2 and 1+1=2 1+2=3 2+3=5 etc. the Golden ratio appears as the limit of succesive terms in the Fibonacci sequence: \[\lim_{n \rightarrow \infty}n/(n+1)\] where n is the nth term in the sequence. In nature it appears everywhere from plants to crystals. It even plays an integral role in the latest Nobel Prize awarded for chemistry: http://www.nobelprize.org/nobel_prizes/chemistry/laureates/2011/sciback_2011.pdf

OpenStudy (turingtest):

sorry n=2 int the sequence is 1+0=1

OpenStudy (anonymous):

the height of a human divided by the height of their navel is very close to golden ratio -believe it or not its true!!)